OpenStudy (anonymous):

what is the answer to this question: y= integral of (e^x^2) dx from 0 to infinity

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

explain how please....

OpenStudy (anonymous):

oh never mind it says from 0 to infinity right?

OpenStudy (anonymous):

yes sir

OpenStudy (zarkon):

\[\infty\]

OpenStudy (anonymous):

please show your steps....thank you

OpenStudy (anonymous):

you do the integration by parts, but there is no reason to because this thing gets huge

OpenStudy (anonymous):

what zarkon said. think of what the graph of \[y=e^x\] looks like

OpenStudy (anonymous):

no what that this integral converges

OpenStudy (anonymous):



OpenStudy (anonymous):

yes i know the graph of y=e^x but i need the full solution to this problem

OpenStudy (zarkon):

\[e^{x^2}\ge1\] for all \[x\ge 0\]

OpenStudy (anonymous):

@zarkon got a better explanation?

OpenStudy (zarkon):

\[\int\limits_{0}^{\infty}e^{x^2}dx\ge\int\limits_{0}^{\infty}1dx=\infty\]

OpenStudy (anonymous):

i already know the answer to this problem from wolfram but need to know the steps to solve it

OpenStudy (anonymous):

oh i even read the problem wrong. i thought it said \[\int_0^{\infty} x^2e^x dx\] i am going to make a bet. my bet is that the question is \[\int\limits_{0}^{\infty}e^{-x^2}dx\]

OpenStudy (zarkon):

that would make for a more interesting problem

OpenStudy (anonymous):

yes sir thats the problem

OpenStudy (anonymous):

this question is from advanced statistics class

OpenStudy (zarkon):

write as a double integral and integrate in polar coordinates

OpenStudy (anonymous):

zarkon...please show your method

OpenStudy (zarkon):

\[I=\int\limits_{0}^{\infty}e^{-x^2}dx\] \[I^2=\int\limits_{0}^{\infty}e^{-x^2}dx\int\limits_{0}^{\infty}e^{-y^2}dy\] \[I^2=\int\limits_{0}^{\infty}\int\limits_{0}^{\infty}e^{-(x^2+y^2)}dxdy\] now switch to polar coordinates

OpenStudy (anonymous):

but that would give me r^2 instead of x^2 and put me back to same problem

OpenStudy (zarkon):

\[I^2=\int\limits_{0}^{\infty}\int\limits_{0}^{\infty}e^{-(x^2+y^2)}dxdy\] \[I^2=\int\limits_{0}^{\pi/2}\int\limits_{0}^{\infty}e^{-r^2}r\,dr\,d\theta\] \[I^2=\int\limits_{0}^{\pi/2}\left[-\frac{1}{2}e^{-r^2}\right]_{0}^{\infty}\,d\theta\] \[I^2=\int\limits_{0}^{\pi/2}\frac{1}{2}\,d\theta=\frac{\pi}{4}\] \[I=\frac{\sqrt{\pi}}{2}\]